OpenStudy (anonymous):

How to add cos(40pi*t+pi)+cos(-40*pi*t-pi)

OpenStudy (anonymous):

I think cos(-40*pi*t-pi) becomes -cos(40*pi*t+pi)?

OpenStudy (anonymous):

oh snap sorry I meant How to add cos(40pi*t-pi)+cos(-40*pi*t-pi)

OpenStudy (zehanz):

No, for every x, cos(-x)=cos(x); cos (x) is an even function.

OpenStudy (anonymous):

still slightly confused

OpenStudy (zehanz):

This does give you a start, btw

OpenStudy (anonymous):

one is a shift of -pi and one is a shift of +pi?

OpenStudy (zehanz):

It means cos(-40pi*t - pi)=cos(40pi*t+pi)

OpenStudy (anonymous):

my question was actually cos(40pi*t-pi) sorry so would cos(-40pi*t-pi) still equal that

OpenStudy (zehanz):

Not really, you have 40pi*t, so to see the shift, you would have to write it as: cos(40pi(t-1/40)), so the shift is only 1/40.

OpenStudy (anonymous):

alright thanks for your help!

OpenStudy (anonymous):

I think I know what I did wrong thank you

OpenStudy (zehanz):

Maybe it would be a good idea to use the formula of Simpson here: \[\cos(x)+\cos(y)=2\cos \frac{ x+y }{ 2 }\cos \frac{ x-y }{ 2 }\]

OpenStudy (zehanz):

YW!